OpenStudy (anonymous):

Assume that sin(x) equals its Maclaurin series for all x. Use the Maclaurin series for sin(8 x^2) to evaluate the integral from 0 to 0.8 sin(8 x^2)dx Your answer will be an infinite series. Use the first two terms to estimate its value

OpenStudy (experimentx):

is it \[ \int_{0}^{0.8} \sin (8x^2) dx = \int_{0}^{0.8} (8x^2/1! - (8x^2)^3/3! + (8x^2)^5/5! -(8x^2)^7/7! + ...) dx\]

OpenStudy (experimentx):

Does not work !!!

OpenStudy (amistre64):

should we wrap it about zero? i can never remember if the mac or the taylow is wrapped at zero

OpenStudy (experimentx):

i think it's mac

OpenStudy (amistre64):

the derivative of sin(8x^2) are: sin(8x^2) 16x cos(8x^2) -16^2 x^2 sin(8x^2) -16^3 x^3 cos(8x^2) which at x=0 simply goes zero all over, so we might have to pick a different center

OpenStudy (amistre64):

x=pi might work better

OpenStudy (amistre64):

or, x = sqrt(pi) lol

OpenStudy (amistre64):

well, the mac wraps about zero according to the google

OpenStudy (amistre64):

the wolfs got a series expansion about zero, but it doesnt say how

OpenStudy (experimentx):

can you post the link??

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sin%288x%5E2%29

OpenStudy (experimentx):

thanks ... the expansion at zero looks like mine ... but i tested integrating first few terms, and it doesn not work at all.